OpenStudy (anonymous):

Please help me do this problem with work shown - find the speed of a 0.150 kg baseball that initially travels east at 40.0 m/s and is then hit with a westward force of 125 N over at 0.010s

OpenStudy (anonymous):

FT=m(delta v) -125(.01)=.15(Vf-40) -8.333=Vf-40 31.667=Vf